On a test of 75 questions, I received a score of 83.
How many questions did I miss?

Answer by richwmiller(17219)

Do all the questions have the same value?
Is there partial credit?
Are you penalized for wrong answers?
assuming all have the same value, no partial credit and no penalty for wrong answers
83/100=x/75
100x=75*83
x=75/100*83
x=3/4*83
x=62.25 which can't be right
75-62.25=12.75 missed
If you missed 12 then 63/75 is clearly 84
If you missed 13 then 62/75 is 82.667 which could be rounded to 83
you missed 13